OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL Sam uses an equation to calculate the total number of apples contained in any number of shipping crates. which of the following is an equation?

OpenStudy (anonymous):

\[A. n = a * c \] \[B. n \ge a *c \] \[C. n \le a*c \] \[D. n \div a*c \]

OpenStudy (anonymous):

Remember than an equation has an equal sign. So that leaves only one answer.

OpenStudy (anonymous):

you say A.

OpenStudy (anonymous):

@mipie13

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

is it A